Operation Summary for September 30, 2009 - October 2, 2009 |
Wednesday Sept. 30 The day shift began with Tevatron (TeV) personnel and operators investigating the TeV abort, Operations monitoring the Recycler stash (175.8E10), with the Antiproton Source stacking (14.5E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store. During the day shift, Operations reported that TeV support personnel replaced the control module and the contactor for a TeV quadrupole power supply (CQ7). Operations reported that a LCW leak on the transformer for the Neutron Therapy Facility (NTF) 58 Degree power supply caused the NTF water skid to trip off. Power supply personnel reported that the transformer was bad, so a replacement supply was moved into position. It will be hooked up on Thursday. Vacuum technicians accessed the TeV sectors A-E to look for an air leak, which then disappeared. They replaced a transducer in the TeV sector E3 stairwell. At 10:30 AM, they reported finding the switch for T:SD to be (thermally) hot. TeV personnel had it changed out by 3 PM. At 3:50 PM, the TeV quenched during turn-on due to a TECAR failure. During the evening shift, operators found the TECAR crate power supply had failed so the backup crate was installed. After operators had replaced the TECAR crate, Operations put the TeV into HEP shot setup at 8:54 PM. Operations established store 7231 at 10:38 PM with an average initial luminosity (AIL) of 255.39E30. Thursday Oct. 1 The midnight shift began with Operations monitoring store 7231 (184.52E30) and the stash (84.1E10), with Pbar stacking (20.6E10), and with NuMI and MiniBooNE taking beam. At 4:13 AM, cryo system technicians reported issues at houses C4 and E1, Operations and CHL monitored the two houses. TeV personnel conducted end-of-store studies from 10 to 10:30 AM, then Operations terminated store 7231. Operations prepared the TeV for an access to check on a leaking Kautzky valve at sector C4. The cryo system techs determined that the valve needed to be changed, so the house was warmed up to 20 Kelvin. Operations reported that the sector C4 cryo system cooldown was hampered by contamination. Cryo system techs began de-riming the system shortly after 8 PM. Friday Oct. 2 The midnight shift began with TeV C4 cryo work in progress, with Operations monitoring the stash (461.5E10), with Pbar stacking (18.4E10), and with NuMI and MiniBooNE taking beam. CHL personnel reported at 6:55 AM that they had completed the C4 cryo decontamination and had recovered the house. Operations began the TeV checkout and turn-on at 7:15 AM. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow Recycler and Pbar studies, Switchyard Septa work, and allow Linac personnel to tune on the new NTF bend magnet power supply. |
|
More Information Comments and Suggestions |
| last modified 10/02/2009 email Fermilab |
| Security, Privacy, Legal | |